A cable suspended between two posts that are the same height and 10 meters apart has a sag of 1 meter. If the cable hangs in the form of a parabola, find its equation, taking the lowest point as the origin.
Answer by stanbon(75887)

A cable suspended between two posts that are the same height and 10 meters apart has a sag of 1 meter. If the cable hangs in the form of a parabola, find its equation, taking the lowest point as the origin.
---------
Sketch the figure on an x/y coordinate system.
---------
Place the lowest point at (0,0)
The top of the posts will be at (-5,10) and (5,10)
-----
Procedure::
Form:: ax^2 + bx + c = y
Using (-5,10) you get: 25a - 5b = 10
Using (5,10) you get: 25a + 5b = 10
-------
50a = 20
a = 2/5
-----
Solve for "b"::
25a + 5b = 10
25a + 2 = 10
25a = 8
a = 8/25
-----
Equation:
y = (8/25)x^2 + (2/5)x
-----------
Cheers,
Stan H.
----------